# LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding

### v1

@@ Abstract
Despite advances in legal NLP, no comprehensive evaluation covering multiple legal-specific LLMs currently exists for contract classification tasks in contract understanding. To address this gap, we present an evaluation of 10 legal-specific LLMs on three English-language contract understanding tasks and compare them with 7 general-purpose LLMs. The results show that legal-specific LLMs consistently outperform general-purpose models, especially on tasks requiring nuanced legal understanding. Legal-BERT and Contracts-BERT establish new SOTAs on two of the three tasks, despite having 69% fewer parameters than the best-performing general-purpose LLM. We also identify CaseLaw-BERT and LexLM as strong additional baselines for contract understanding. Our results provide a holistic evaluation of legal-specific LLMs and will facilitate the development of more accurate contract understanding systems.
LLMs for Law: Evaluating Legal-Specific LLMs on Contract Understanding
Amrita Singhthanks: These authors contributed equally to this work., H. Suhan Karaca11footnotemark: 1, Aditya Joshi, Hye-young Paik, Jiaojiao Jiang School of Computer Science and Engineering University of New South Wales (UNSW), Sydney
@@ 1 Introduction
Recent work suggests that open-source legal-specific LLMs offer a promising, cost-effective, and privacy-preserving alternative to general-purpose LLMs Bhambhoria et al. (2024); Chalkidis et al. (2020).
Prior Work / Ours
Legal-Specific LLMs UNFAIR-ToS LEDGAR LEXDEMOD
Legal-BERT ✓/✓ ✓/✓ ✕/✓
Contracts-BERT ✕/✓ ✕/✓ ✓/✓
Legal-RoBERTa ✕/✓ ✕/✓ ✕/✓
CaseLaw-BERT ✓/✓ ✓/✓ ✕/✓
PoL-BERT ✕/✓ ✕/✓ ✕/✓
InLegalBERT ✕/✓ ✕/✓ ✕/✓
InCaseLawBERT ✕/✓ ✕/✓ ✕/✓
CustomInLawBERT ✕/✓ ✕/✓ ✕/✓
LexLM ✕/✓ ✕/✓ ✕/✓
Legal-XLM-R ✕/✓ ✕/✓ ✕/✓
Table 1: Comparison of our legal-specific LLMs evaluation and coverage with prior work across three contract datasets (tasks): ‘UNFAIR-ToS’ (Unfair Contractual Terms Identification, Lippi et al. (2019); Chalkidis et al. (2022)), ‘LEDGAR’ (Contract Provision Topic Classification, Tuggener et al. (2020); Chalkidis et al. (2022)), and ‘LEXDEMOD’ (Agent-Specific Deontic Modality Detection, Sancheti et al. (2022)). Terms in inverted commas refer to dataset names, tasks are in parentheses. ✓= model inclusion, ✕= model exclusion.
@@
However, despite their advantages, these models remain significantly underutilized in current legal NLP downstream tasks. As illustrated in Table 1, legal-specific LLMs are rarely evaluated in prior work on three popular and freely available contract understanding tasks such as Unfair Contractual Terms Identification Lippi et al. (2019); Chalkidis et al. (2022), Contractual Provision Topic Classification Tuggener et al. (2020); Chalkidis et al. (2022), and Agent-Specific Deontic Modality Detection Sancheti et al. (2022). Despite the legal nature of documents/tasks, researchers have continued to favor general-purpose LLMs over legal-specific LLMs. In some cases, legal-specific LLMs are excluded entirely. For instance, recent studies such as Guha et al. (2023) and Singh et al. (2024), which explicitly focus on legal downstream tasks, do not include any legal-specific LLMs in their benchmarking evaluations. Therefore, this paper addresses the Research Question (RQ): How do legal-specific LLMs perform compared to general-purpose LLMs on nuanced legal tasks like contract understanding?
Dataset
Contract
Type
Task Task Type
Train/Dev/
Test Instances
Classes
UNFAIR-ToS
Chalkidis et al. (2022)
Terms of
Service
(Consumer
Contract)
Unfair Contractual
Terms Identification
Multi-label
Classification
5,532/2,275/
1,607
9
LEDGAR
Chalkidis et al. (2022)
Exhibit-10
Material
Contract
Contract Provision
Topic Classification
Multi-class
Classification
60,000/10,000/
10,000
100
LEXDEMOD
Sancheti et al. (2022)
Lease
Contract
Agent-Specific Deontic
Modality Detection
Multi-label
Classification
4,282/330/
1,777
7
Table 2: Overview of Datasets used for Benchmarking Legal-specific LLMs.
@@
To address this question, we present a comprehensive evaluation of 10 open-source legal-specific LLMs with 7 general-purpose LLMs across the three distinct contract understanding tasks. Our results reveal consistent improvements in performance for legal-specific LLMs, particularly on tasks where legal and domain-specific semantics are critical. This benchmark serves as a resource for the community, offering a clearer understanding of model suitability and performance across tasks and model types. The contributions of this work are as follows: (a) To the best of our knowledge, we present the first benchmarking of multiple legal-specific LLMs across multiple contract understanding tasks; (b) We systematically compare their performance with that of general-purpose LLMs; (c) We identify model strengths, weaknesses, and task-specific challenges, offering insights for future research and deployment.
@@ 2 Contract Classification Tasks and Datasets
2.1 Dataset Selection Desiderata
Based on following factors, we select the legal contract datasets and tasks:
Language: English-language contract datasets are selected due to their availability, provide consistent benchmarking for future legal-specific models in the global research community, enable comparison with past benchmarked models.
Relevance and Diversity: The focus is on contract classification tasks that reflect real-world contract review and analysis challenges, and that test a model’s understanding of legal language, structure, and semantics. As shown in Table 2, three distinct tasks are selected, each using a different dataset and representing a unique contract classification scenario in terms of dataset size and number of classes.
Difficulty: Datasets are chosen where SOTA general-purpose language models do not achieve near-perfect performance Lippi et al. (2019); Tuggener et al. (2020); Sancheti et al. (2022), ensuring that benchmarking legal-specific language models remains challenging.
Availability & Size: Public, well-documented datasets are used, each large enough for stable training and evaluation. Proprietary, non-public, and very small datasets (under 3K sentences) are avoided to ensure reproducibility and generalizability. This criterion modifies and adapts the selection guidelines of Chalkidis et al. (2022).

### v2

# Evaluating Customized vs. Generalist Transformer-based Models for Legal Contract Classification

@@ Abstract
Despite advances in legal NLP, no comprehensive evaluation of Transformer-based models customized for legal tasks (referred to as ‘legal-specific’ models in this paper) exists for contract classification tasks. To address this gap, we present an evaluation of 13 legal-specific transformer-based models on 3 English-language contract classification tasks and compare them with 9 generalist models. The results show that legal-specific models consistently outperform generalist models, especially on tasks requiring nuanced legal understanding. They also help reduce misclassification of rare classes in imbalanced datasets. Legal-BERT and Contracts-BERT establish new SOTAs on two of the three tasks, despite having 69% fewer parameters than the best-performing generalist models. We also identify CaseLaw-BERT and LexLM as strong additional baselines for contract classification. Our results highlight the shortcomings of generalist models, emphasizing the need for domain-specific customization, particularly in the context of legal applications.
Evaluating Customized vs. Generalist Transformer-based Models for Legal Contract Classification
Amrita Singh, H. Suhan Karaca, Aditya Joshi, Hye-young Paik, Jiaojiao Jiang School of Computer Science and Engineering University of New South Wales (UNSW), Sydney 11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Recent work suggests that open-source legal-specific models offer a promising, cost-effective, and privacy-preserving alternative to generalist models Singh et al. (2025); Bhambhoria et al. (2024); Chalkidis et al. (2020). However, despite their advantages, these models remain significantly underutilized in current legal contract classification tasks, as noted by Singh et al. (2025). As illustrated in Figure 1, legal-specific models are rarely evaluated, and mostly generalist models are preferred in prior work on three popular and freely available contract classification tasks: Unfair Contractual Terms Identification Lippi et al. (2019); Chalkidis et al. (2022), Contractual Provision Topic Classification Tuggener et al. (2020); Chalkidis et al. (2022), and Agent-Specific Deontic Modality Detection Sancheti et al. (2022).
Figure 1: Comparison of our legal-specific models evaluation and coverage with prior work.
@@
Despite the legal nature of documents/tasks, researchers have continued to favor generalist models over legal-specific models. This holds true not only for decoder models (which is the predominant context in which ‘customization’ is used) but also for encoder models that must be adopted to the legal domain. In some cases, legal-specific models are excluded entirely. For instance, recent studies such as Guha et al. (2023) and Singh et al. (2024), which explicitly focus on legal downstream tasks, do not include any legal-specific models in their benchmarking evaluations. Therefore, this paper addresses the Research Question (RQ): How do legal-specific models perform compared to generalist models on legal contract classification tasks? To address this question, we present a comprehensive evaluation of 13 open-source legal-specific models with 9 generalist models across the three distinct contract classification tasks. Our results reveal consistent improvements in performance for legal-specific models, particularly on tasks where legal and domain-specific semantics are critical. This benchmark serves as a resource for the community, offering a clearer understanding of model suitability and performance across tasks and model types. The contributions of this work are as follows: (a) To the best of our knowledge, we present the first benchmarking of multiple legal-specific models across multiple contract classification tasks; (b) We systematically compare their performance with that of generalist models; (c) We identify model strengths, weaknesses, and task-specific challenges, offering insights for future research and deployment.
@@ 2 Contract Classification Tasks and Datasets
To evaluate the effectiveness of models customized for the legal setting vis-a-vis generalist models, we consider the following factors to select the datasets and tasks. We select the language to be English due to the availability of datasets and models. Publicly available, well-documented datasets are used, each large enough for stable training and evaluation. Proprietary, non-public, and very small datasets (under 3K sentences) are avoided to ensure reproducibility and generalizability. This criterion modifies and adapts the selection guidelines of Chalkidis et al. (2022). Datasets are chosen where SOTA generalist language models do not achieve near-perfect performance Lippi et al. (2019); Tuggener et al. (2020); Sancheti et al. (2022), ensuring that benchmarking legal-specific language models remains challenging. For the chosen tasks to reflect relevance and diversity, we select tasks that test a model’s understanding of legal language, structure, and semantics. As shown below, three distinct tasks are selected, each using a different dataset and representing a unique contract classification scenario in terms of dataset size and number of classes.
Consequently, the following datasets are selected: UNFAIR-ToS: The UNFAIR-ToS dataset (train (5.5k), dev (2.3k), and test (1.6k)) from Chalkidis et al. (2022) is a multi-label classification dataset that is used to identify unfair contractual terms in Terms of Service (ToS) documents from online platforms like YouTube. Each sentence is annotated with one or more of 8 unfairness categories, plus 1 unlabeled class for sentences that do not indicate any potential violation of European consumer law EU consumer protection law .
LEDGAR: The LEDGAR dataset (train (60k), dev (10k), and test (10k)) from Chalkidis et al. (2022) is used to classify the principal topic of provisions in Exhibit 10 material contracts (e.g., employment, lease, non-disclosure) filed with the US Securities and Exchange Commission (SEC) via EDGAR . Each provision (paragraph) is labeled with one of 100 contract topics, making it a multi-class classification task.
LEXDEMOD: The LEXDEMOD dataset (train (4.2k), dev (330), and test (1.7k)) from Sancheti et al. (2022) detects deontic modality in agent-based contract clauses from lease agreements sourced from the LEDGAR dataset. Each clause (sentence) is annotated with one or more of 6 deontic modality types plus 1 none class , making it a multi-label classification task. Labels are linked to an agent (party) in the sentence, representing their deontic status (e.g., Obligation, Entitlement, Prohibition). The train/dev/test split is as reported in the original paper. Statistics and examples for all datasets are in Appendix A while the class distribution is in Figure 3.
UNFAIR-ToS LEDGAR LEXDEMOD
Method Model # Params $\mu$-F1 m-F1 $\mu$-F1 m-F1 $\mu$-F1 m-F1
BERT 110M 95.6 81.3 87.6 81.8 - 75.61
RoBERTa-base 125M 95.2 79.2 87.9 82.3 - 75.66
DeBERTa 139M 95.5 80.3 88.2 83.1 - -
Longformer 149M 95.5 80.9 88.2 83.0 - -
BigBird 127M 95.7 81.3 87.8 82.6 - -
Baselines reported from: Chalkidis et al. (2022), Sancheti et al. (2022) Generalist Models RoBERTa-large 355M 95.8 81.6 88.6 83.6 - 77.88
Biswas et al. (2025) Legal-KDD-Student DistilBERT 66M 95.9 76.2 88.2 82.3 - -
Shin and Kaneko (2025) GATs Hierarchical BERT 110M 95.6 81.5 88.4 83.1 - -
Llama-3.2 3B 95.9 80.3 85 76.1 76.2 71.4
Generalist Models Mistral 7B 96.0 80.7 86.4 79.2 76.0 71.2
Legal-BERT 110M 96.0 82.2 88.2 82.5 81.23 78.01
Contracts-BERT 110M 96.2 83.4 87.9 82.2 80.17 77.71
Legal-RoBERTa 125M 95.4 81.1 87.7 81.9 80.12 76.70
CaseLawBERT 110M 96.1 83.2 87.6 80.9 80.32 77.75
PoL-BERT 340M 94.6 77.9 86.0 79.1 41.35 15.75
InLegalBERT 110M 95.6 81.7 87.9 82.0 80.21 77.89
InCaseLawBERT 110M 95.5 81.1 87.5 82.1 79.16 76.83
CustomInLawBERT 110M 95.5 79.9 87.7 81.8 78.16 75.35
LexLM 124M 95.9 81.7 87.8 81.3 80.39 77.46
Legal-XLM-R 184M 94.9 78.2 87.7 81.7 80.62 77.56
LexT5 220M 95.4 79.8 84.9 76.1 76.50 73.30
AdaptLLM 7B 96.5 83.2 85.2 76.8 76.2 70.0
Proposed Legal-specific Models SaulLM 7B 96.0 81.0 86.6 79.4 76.5 72.7
Table 1: Performance of legal-specific and generalist models on three tasks: UNFAIR-ToS, LEDGAR, LEXDEMOD. Metrics: micro-F1 ($\mu$-F1) and macro-F1 (m-F1). Blue highlights the best generalist, and red highlights the best legal-specific model performance.
Mean $\pm$ Std
Legal Specific Models $\mu$-F1 m-F1
Legal-BERT 88.48 ± 6.03 80.90 ± 2.05
Contracts-BERT 88.09 ± 6.55 81.10 ± 2.45
Legal-RoBERTa 87.74 ± 6.24 79.90 ± 2.29
CaseLawBERT 88.01 ± 6.45 80.62 ± 2.23
PoL-BERT 73.98 ± 23.34 57.58 ± 29.58
InLegalBERT 87.90 ± 6.28 80.53 ± 1.87
InCaseLawBERT 87.39 ± 6.67 80.01 ± 2.29
CustomInLawBERT 87.12 ± 7.09 79.02 ± 2.71
LexLM 88.03 ± 6.33 80.15 ± 1.91
Legal-XLM-R 87.74 ± 5.83 79.15 ± 1.82
LexT5 85.60 ± 7.73 76.40 ± 2.66
AdaptLLM 85.97 ± 8.31 76.67 ± 5.39
SaulLM 86.37 ± 7.96 77.70 ± 3.60
Table 2: Aggregated scores (Mean $\pm$ Std) across three contract classification tasks. Red, blue, and green highlights indicate the first, second, and third best performances, respectively.
3 Experiment Setup
We perform task-specific fine-tuning using 13 legal-specific models on three datasets: LEDGAR, UNFAIR-ToS, and LEXDEMOD. We consider ten pre-trained encoder-based legal-specific models for fine-tuning. Nine of these are base-variant encoder models: Legal-BERT Chalkidis et al. (2020), Contracts-BERT Chalkidis et al. (2020), LegalRoBERTa Geng et al. (2021), CaseLaw-BERT Zheng et al. (2021), InLegalBERT, InCaseLawBERT, and CustomInLawBERT Paul et al. (2023), Legal-XLM-R Niklaus et al. (2024), and LexLM Chalkidis* et al. (2023). One large-variant model, PoL-BERT Henderson et al. (2022), is included, as its base version is not present. We also evaluate two legal decoder-based models, AdaptLLM Cheng et al. (2024) and SaulLM-7B Colombo et al. (2024), along with one legal encoder-decoder model, LexT5 T.y.s.s et al. (2024), as it is the only model of its kind among legal-specific decoder-based and encoder-decoder models. Table 5 in Appendix B summarizes the key characteristics of each model, and a detailed description of them is provided in the same appendix. A detailed experimental setup is provided in Appendix C. We compare these 13 legal-specific models with 9 generalist models. These include six base variant encoder models: BERT Devlin et al. (2019), RoBERTa-base Liu et al. (2019), DeBERTa He et al. (2021), Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), and DistilBERT Sanh (2019), along with one large variant, RoBERTa-large Liu et al. (2019). We also compare two generalist decoder-based models, Llama-3.2 Grattafiori et al. (2025) and Mistral Jiang et al. (2023).
4 Results and Analysis
Generalist models: Table 1 reports the test results of models across all three tasks, comparing legal-specific and generalist models. Among generalist models, RoBERTa-large (355M) outperforms others, including larger decoder-based models like Llama-3.2 (3B) and Mistral (7B). Figure 3 shows that all datasets suffer from extreme long-tail class imbalance. Table 1 confirms that decoder-based models perform poorly on rare classes, as reflected in their lower macro-F1 scores, while RoBERTa-large performs consistently better. This suggests that encoder-based models are more robust to extreme class imbalance than decoder-based models. Despite having fewer parameters, RoBERTa-large achieves higher micro-F1 and macro-F1, demonstrating that bidirectional context learning and discriminative fine-tuning are more crucial than model scale for legal classification tasks with long-tail distributions.
Legal-specific models: Legal-specific models such as Contracts-BERT and Legal-BERT (110M) outperform RoBERTa-large on UNFAIR-ToS and LEXDEMOD, respectively, despite having 69% fewer parameters. RoBERTa-large remains the best model for LEDGAR. Still, Legal-BERT delivers equivalent performance compared to generalist base variant models on this task, suggesting that both model size and task characteristics influence performance. The larger, legal-specific encoder-based models may be better suited for LEDGAR, as the dataset is very large. Overall, legal-specific base models deliver competitive performance and set new SOTAs on two of the three tasks, demonstrating the effectiveness of domain-specific pretraining, even at the base variant of models.
Generalist models vs. Legal-specific models: We further analyze the errors made by the best generalist model (RoBERTa-large) and examine whether the best legal-specific model (Contracts-BERT) corrects them on the Unfair-ToS dataset. As shown in Figure 2, RoBERTa-large majorly misclassifies clauses belonging to rare categories, particularly Limitation of Liability and Unilateral Termination, whereas Contracts-BERT correctly classifies them, suggesting that domain-specific pretraining on contract corpora (refer Table 5) enables better representation of rare legal clause types that generalist models fail to capture. This analysis is further extended to decoder-based models by comparing generalist and their corresponding legal-specific variants on the LEDGAR dataset, as shown in Figure 2. Specifically, Mistral and its legal-specific counterpart SaulLM, both sharing the same base architecture, are compared alongside Llama-3.2 and AdaptLLM, where AdaptLLM is built on Llama-1 but shares the same Llama architecture lineage. Across all cases, the results demonstrate that domain-specific pretraining helps in reducing the misclassification of rare classes in long-tail legal text classification.
Figure 2: Example outputs showing customized models (green squares) correctly predict examples which were mis-classified by generalist models (red squares).
Legal-specific models as baselines: Table 2 presents aggregated test scores (Mean $\pm$ Std) across the three contract classification tasks. Despite class imbalance in all tasks, Legal-BERT achieves the highest aggregated $\mu$-F1, while Contracts-BERT leads in m-F1. Across both metrics, the top positions are consistently held by four legal-specific models: Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM. We conclude that these four models, Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM, should be considered strong baselines for contract classification tasks.
Limitations of Recent Legal-Specific Models: Several recent legal-specific models, such as PoL-BERT, CustomInLawBERT, LexLM, Legal-XLM-R are pre-trained on large-scale legal corpora. Models like InLegalBERT and InCaseLawBERT are built on legal-specific base models rather than generalist models. However, older models like LegalBERT and ContractsBERT, pre-trained on just 354k and 76k legal documents (Table 5), still outperform recent base-variant legal-specific models (Table 2). This counterintuitive result can be attributed to a data distribution mismatch: more data only improves performance when it is in-distribution with the task. Recent models are trained on a broad mixture of legal genres, including court cases, legislation, and patents, which dilutes the contract-specific signal. Consequently, contracts remain underrepresented relative to other genres in these corpora. We conclude that future legal-specific models should incorporate a more diverse and representative set of contract documents, balanced alongside other legal genres, to improve performance on contract-based downstream tasks.
5 Conclusion
This study benchmarks 13 legal-specific and 9 generalist models across three contract classification tasks. Encoder-based models outperform larger decoder-based models in long-tail legal classification, with domain-specific pretraining consistently improving rare class recognition. Legal-specific models set new state-of-the-art results on two of the three tasks, despite having 69% fewer parameters than the best-performing generalist model. Legal-BERT, Contracts-BERT, CaseLaw-BERT, and LexLM serve as strong baselines for future research. Notably, older legal-specific models trained on smaller, more focused corpora outperform recent models trained on larger, genre-diverse corpora, highlighting that in-distribution pretraining is more critical than scale. Future work will explore retrieval-augmented approaches with legal-specific models to further enhance performance on rare and underrepresented legal clause types.
Limitations
The limited availability of contract benchmark datasets in languages other than English presents a challenge for multilingual extension. As a result, this study focuses solely on English-language contract tasks, with evaluation on non-English data left for future research. This work also concentrates on the nuances of contract language and does not assess performance on other legal text types, such as statutes, court decisions, or legal opinions. Future research should broaden this evaluation to encompass a wider range of legal genres, recognizing that no single study can fully capture the entire legal domain. Additionally, this paper focuses on domain generalization alone while selecting the legal domain.
Ethical Considerations
This study uses only publicly available datasets, LEDGAR, UNFAIR-ToS, and LEXDEMOD, all of which contain contract clauses from publicly available contract documents. LEDGAR is derived from public U.S. SEC EDGAR filings, UNFAIR-ToS from company Terms of Service, and LEXDEMOD from lease clauses sourced from LEDGAR. This research does not offer legal advice, predict individual outcomes, or automate decisions affecting rights. It focuses solely on evaluating the performance of legal-specific models to inform future tools and research. While these models can support legal professionals, they are not substitutes for legal expertise. We acknowledge potential ethical risks if outputs are misused or inaccurate. By open-sourcing our evaluations, we aim to reduce reliance on proprietary tools, promote transparency, and expand access to legal AI research and development.
Appendix A Dataset Statistics and Illustrative Examples
Dataset Contract Type Task Task Type Train /Dev/ Test Class
UNFAIR-ToS Chalkidis et al. (2022) Terms of Service (Consumer Contract) Unfair Contractual Terms Identification Multi-label Classification 5,532/2,275/ 1,607 9
LEDGAR Chalkidis et al. (2022) Exhibit-10 Material Contract Contract Provision Topic Classification Multi-class Classification 60,000/10,000/ 10,000 100
LEXDEMOD Sancheti et al. (2022) Lease Contract Agent-Specific Deontic Modality Detection Multi-label Classification 4,282/330/ 1,777 7
Table 3: Overview of Datasets used for Benchmarking Legal-specific Models.
This appendix provides an overview of the datasets used for benchmarking legal-specific models in Table 3 and presents labeled examples from all three datasets to aid understanding. The class distribution of all three datasets is given in Figure 3. We provide our own rationales explaining label types, which the datasets do not explicitly include. These explanations clarify why specific labels apply to given clauses (sentences) or provisions (paragraphs), as detailed in Table 4. Legal contract classification involves longer texts than typical NLP tasks like tweets or reviews. Legal-specific Transformer models such as Legal-BERT process up to 512 sub-word tokens, but many LEDGAR paragraphs exceed this limit. Figure 4 shows that numerous LEDGAR paragraphs surpass the standard context window, requiring truncation or other methods to handle long inputs. Additionally, legal texts contain specialized terminology (legalese), increasing classification complexity compared to general text.
Figure 3: Class Distribution across all three datasets
Dataset Example Label Rationale for Assigned Labels (Provided by us for better understanding)
This Amendment may be executed by one or more of the parties hereto on any number of separate counterparts, and all of said counterparts taken together shall be deemed to constitute one and the same instrument. This Amendment may be delivered by facsimile or other electronic transmission of the relevant signature pages hereof. Counterparts This sentence states that the Amendment may be executed in multiple counterparts and that together they form a single agreement, which is a standard counterparts clause used to validate separately signed copies as one binding document.
THIS AMENDMENT SHALL BE GOVERNED BY, AND INTERPRETED IN ACCORDANCE WITH, THE LAW OF THE STATE OF NEW YORK . The other provisions of Article IX of the Credit Agreement shall apply to this Amendment to the same extent as if fully set forth herein. Governing Laws This sentence specifies that New York law will govern and interpret the Amendment, which is a standard governing law clause that establishes the legal jurisdiction and framework for resolving disputes.
LEDGAR Sublessee leases the Aircraft in its “as is, where is” condition. The only services, rights, or warranties to which the Sublessee is entitled to under this Sublease are those to which the Sublessor is provided under the Prime Lease. Warranties The sentence is labeled as warranties because it defines the rights and guarantees the Sublessee receives and limits those warranties to what the Sublessor has under the Prime Lease.
Niantic further reserves the right to remove any User Content from the Service at any time and without notice and for any reason. Content removal This sentence is labeled as content removal unfair contractual term because it gives the provider full control to remove content at any time, for any reason, and without notice.
amazon reserves the right to refuse service, terminate accounts , terminate your rights to use amazon services, remove or edit content , or cancel orders in its sole discretion. Unilateral termination, Content removal This sentence is labeled as unilateral termination and content removal because it allows Amazon to end services, remove content, or cancel orders at its sole discretion, without notice, creating an imbalance of power.
UNFAIR-ToS Outside the United States and Canada. If you acquired the if you acquired the application in any other country, the laws of that country apply. None The sentence is labeled as none because it does not belong to any of the unfair contractual term types and is actually a fair clause.
[lessee] Lessee will not create or permit to be created or to remain , and will promptly discharge , any lien , encumbrance or charge (including without limitation any mechanic ’s , laborer ’s or materialman ’s lien ) against the Premises or any part thereof arising from Lessee ’s actions. Prohibition, Obligation This sentence imposes a prohibition by forbidding the Lessee (the agent) from creating or allowing liens. It also imposes an obligation by requiring the Lessee to promptly remove any such liens.
[tenant] Tenant may, without Landlord’s consent, before delinquency occurs, contest any such taxes related to the Personal Property. Permission This sentence grants permission to the Tenant (the agent) to contest taxes without needing the Landlord’s consent, as long as it’s done before delinquency.
LEXDEMOD [landlord] Tenant shall promptly notify Landlord of any alleged defaults under the CC&Rs and/or the Oil and Gas Lease . Entitlement The Landlord, as the agent, holds an entitlement to receive notice from the Tenant about alleged defaults.
Appendix B Description of Legal-specific Models